(2, 2) 型の Abel (拡大) 体の類数および類群に関する文献 (2012 年 1 月更新版)

[1] G. L. Dirichlet : Untersuchungen über die Theorie der complexen Zahlen, Bericht Verhand. Königl. Preuss. Akad. Wiss. Jahrg. 1841, 190–194; [jpg], Akademiebibliothek; reprinted in G. Lejeune Dirichlet’s Werke. Herausgegeben auf Veranlassung der Königlich Preussischen Akademie der Wissenschaften von L. Kronecker, Erster Band mit G. Lejeune Dirichlet’s Bildniss, G. Keimer, Berlin, 1889, XXX, S. 503–508; [PDF, JPEG, txt], Gallica; [PDF], UMDL; reprinted by Chelsea Publishing Co., Bronx, New York, 1969; J. Reine Angew. Math. XXII (1841), Hft. 4, 375–378; doi: 10.1515/crll.1841.22.375 [PDF (681 KB)];

PPN243919689_0022_log18.pdf (515,09 KB), GDZ; [PDF],

DigiZeitschriften.

[2] G. L. Dirichlet : Untersuchungen über die Theorie der complexen Zahlen, Abh. Königl. Preuss. Akad. Wiss. Physik.- math. Kl. 1841, 141–161; [jpg], Akademiebibliothek; reprinted in G. Lejeune Dirichlet’s Werke. Herausgegeben auf Veranlassung der Königlich Preussischen Akademie der Wissenschaften von L. Kronecker, Erster Band mit G. Lejeune Dirichlet’s Bildniss, G. Keimer, Berlin, 1889, XXXI, S. 509– 532; [PDF, JPEG, txt], Gallica; [PDF], UMDL; reprinted by Chelsea Publishing Co., Bronx, New York, 1969.

[3] G. L. Dirichlet : Recherches sur les formes quadratiques à coefficients et à indéterminées complexes, J. Reine Angew. Math. XXIV (1842), Heft 4, 291– 371; doi: 10.1515/crll.1842.24.291 [PDF (9310 KB)];

PPN243919689_0024_log32.pdf (7,94 MB), GDZ; [PDF],

DigiZeitschriften; reprinted in G. Lejeune Dirichlet’s Werke. Herausgegeben auf Veranlassung der Königlich Preussischen Akademie der Wissenschaften von L. Kronecker, Erster Band mit G. Lejeune Dirichlet’s Bildniss, G. Keimer, Berlin, 1889, XXXII, S. 533–618; [PDF, JPEG, txt], Gallica; [PDF], UMDL; reprinted by Chelsea Publishing Co., Bronx, New York, 1969.

[4] R. I. Bachmann : Zur Theorie der complexen Zahlen, J. Reine Angew. Math. LXVII (1867), Heft 2, 200– 204; doi: 10.1515/crll.1867.67.200 [PDF (682 KB)];

PPN243919689_0067_log16.pdf (491,37 KB), GDZ; [PDF],

DigiZeitschriften; ERAM 067.1746cj.

[5] D. Hilbert : Ueber den Dirichlet’schen biquadratischen Zahlkörper, Math. Ann. XLV (1894), Heft 3, 309–340; DOI: 10.1007/BF01446682 [xx1vv3j175676513.pdf (1.6 MB)];

PPN235181684_0045_log29.pdf (1,99 MB), GDZ; [PDF], Di-

giZeitschriften; reprinted in David Hilbert Gesammelte Abhandlungen, Erster Band: Zahlentheorie, J. Springer, Berlin, 1932, 5, S. 24–52; PPN237821559_log10.pdf (3,21 MB), GDZ; reprint by Chelsea Publshing Company, Bronx, New York, 1965; Zweite Auflage: Springer-Verlag, Berlin and New York, 1970; JFM 25.0303.01 (reviewed by Minkowski) [html, DVI, ps, PDF, TeX].

[6] G. Herglotz : Über einen Dirichlet’schen Satz, Math. Z. XII (1922), 255–261; DOI: 10.1007/ BF01482079 [q6206705m568475k.pdf (294.9 KB)];

PPN266833020_0012_log20.pdf (538,26 KB), GDZ; [PDF],

DigiZeitschriften; JFM 48.0170.01 (reviewed by Hecke) [html, DVI, ps, PDF, TeX].

[7] S. Kuroda : Über den Dirichletschen Körper, J. Fac. Imp. Univ. Tokyo Sect. I [東京帝國大學理學部紀要. 第 1 類] 4 (1943), 383–406; MR0021031 (9, 12f) (reviewed by M. Krasner; reprinted in Reviews in Number Theory, volume 5A, R14-200) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0061.05901 (reviewed by Z. Suetuna) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[8] H. Hasse : Über die Klassenzahl abelscher Zahlkörper, Mathematische Lehrbucher und Monographien, Band I, Akademie-Verlag, Berlin, 1952, xii+190 pp.; MR0049239 (14, 141a) (reprinted in Reviews in Number Theory, volume 5, R02-10) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0046.26003 (reviewed by S. Kuroda) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; Nachdruck der ersten Aufgabe mit einer Einleitung von J. Martinet : Springer-Verlag, New York and Berlin, 1985, xxii+xii+190 pp.; MR0842666 (87j:11122a) (reviewed by Editors) [html, PDF, DVI, PostScript, BibTeX]; Mathematische Lehrbucher und Monographien, II. Abteilung: Mathematische Monographien, Band 1, Akademie-Verlag, Berlin, 1985; MR0834791 (87j:11122b) (reviewed by Editors) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0668.12004 [html, PDF, XML, ASCII, DVI, PS, BibTeX], 0668.12003 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, Bib-TeX].

[9] T. Kubota : Über die Beziehung der Klassenzahlen der Unterkörper des bizyklischen biquadratischen Zahlkörpers, Nagoya Math. J. 6 (1953), 119–127; euclid.nmj/1118799485 [PDF (757 KB)]; MR0059960 (15, 605e) (reviewed by R. Brauer; reprinted in Reviews in Number Theory, volume 5, R16-12) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0053.21902 (reviewed by S. Kuroda) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[10] T. Kubota : Über den bizyklischen biquadratischer Zahlkörper, Nagoya Math. J. 10 (1956), 65–85; euclid.nmj/1118799770 [PDF (1406 KB)]; MR0083009 (18, 643e) (reviewed by G. Whaples; reprinted in Reviews in Number Theory, volume 5, R16-21) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0074.03001 (reviewed by H. W. Leopoldt) [scanned image].

[11] H. Wada : On the class number and the unit group of certain algebraic number fields, J. Fac. Sci. Univ. Tokyo Sect. IA Math. [東京大学理学部紀要第 I 類 A 数学] 13 (1966), No. 2, 201–209; jfs130209.pdf (308 Kb), URI: http://hdl.handle.net/2261/6067; MR0214565 (35 #5414) (reviewed by T. Kubota; reprinted in Reviews in Number Theory, volume 5, R26-35) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0158.30103 (reviewed by A. Yokoyama) [scanned image].

[12] R. B. Lakein : A Gauss bound for a class of biquadratic fields, J. Number Theory 1 (1969), no. 1, 108–112; doi:10.1016/0022-314X(69)90028-6 [PDF (218 K)]; MR0240073 (39 #1427) (reviewed by D. Pumplün; reprinted in Reviews in Number Theory, volume 5, R16-48) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0208.06103 (reviewed by C.-O. Selenius) [scanned image].

[13] H. Reichardt : Über die Idealklassengruppe des Dirichletschen biquadratischen Zahlenkörpers, Acta Arith. XXI (1972), 323– 327; aa21135.pdf, icm; MR0376615 (51 #12790) (reviewed by K. Lakkis; reprinted in Reviews in Number Theory, volume 5A, R16-234) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0215.07201 (Autorreferat) [scanned image].

[14] P. Kaplan : Divisibilité par 8 du nombre des classes des corps quadratiques dont le 2-groupe des classes est cyclique, et reciprocate biquadratique, J. Math. Soc. Japan [日本数学会欧文誌] 25 (1973), No. 4, 596–608; doi:10.2969/jmsj/02540596, euclid.jmsj/1240435459 [PDF (1229 KB)]; JOI:JST.Journalarchive/jmath1948/25.596 [25_596.pdf (1027 K)], Journal@rchive, JST; MR0323757 (48 #2113) (reviewed by H. Cohn; reprinted in Reviews in Number Theory, volume 5A, R14-224) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0276.12006 (reviewed by R. Bölling) [scanned image].

[15] H. Lang : Über die Klassenzahl eines imaginären bizyklischen biquadratischen Zahlkörpers und seines reell-quadratischern Teilkörpers. I, J. Reine Angew. Math. 262/263 (1973), 18–40; doi: 10.1515/crll.1973.262-263.18 [PDF (2676

KB)]; PPN243919689_0262_0263_log6.pdf (2,30 MB), GDZ;

[PDF], DigiZeitschriften; MR0325573 (48 #3920) (reviewed by K. Lakkis; reprinted in Reviews in Number Theory, volume 5A, R26-211) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0285.12013 (reviewed by A. Yokoyama) [scanned image].

[16] H. Lang : Über die Klassenzahl eines imaginären bizyklischen biquadratischen Zahlkörpers und seines reell-quadratischern Teilkörpers. II, J. Reine angew. Math. 267 (1974), 175– 178; doi: 10.1515/crll.1974.267.175 [PDF (615 KB)];

PPN243919689_0267_log16.pdf (384,50 KB), GDZ; [PDF],

DigiZeitschriften; MR0342492 (49 #7238) (reviewed by K. Lakkis; reprinted in Reviews in Number Theory, volume 5A, R26-211E) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0406.12004 (reviewed by K. Lakkis) [scanned image].

[17] H. Lang und R. Schertz : Bemerkungen über die Klassenzahl eines imaginären bizyklischen biquadratischen Zahlkörpers, Abh. Math. Sem. Univ. Hamburg 42 (1974), 212–216; DOI: 10.1007/BF02993548 [v7386407578n6757.pdf (210.4 KB)]; MR0412149 (54 #276) (reviewed by T. Metsänkylä; reprinted in Reviews in Number Theory, volume 5A, R16-252) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0357.12007 (reviewed by T. Lepistö) [scanned image].

[18] R. A. Mollin : Lower bounds for class numbers of real quadratic and biquadratic fields, Proc. Amer. Math. Soc. 101 (1987), no. 3, 439–444; DOI: 10.1090/S0002-9939-1987-0908645-0 [PDF]; doi:10.2307/2046385 [PDF], JSTOR; MR0908645 (89a:11117) (reviewed by H. Yokoi; reprinted in Reviews in Number Theory, volume 5B, R29-736) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0632.12006 (reviewed by H. Yokoi) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[19] Y. Hayashi : Reelle biquadratische Zahlkörper mit ungerader Klassenzahl, Manuscripta Math. 62 (1988), fasc. 1, 65–82; DOI: 10.1007/ BF01258268 [p18752p163v04w41.pdf

(590.9 KB)]; PPN365956996_0062_log7.pdf (1010,72 KB),

GDZ; [PDF], DigiZeitschriften; MR0958254 (90a:11123) (reviewed by I. Sh. Slavutskiĭ; reprinted in Reviews in Number Theory, volume 5B, R16-662) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0662.12007 (reviewed by E. Brown) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[20] R. I. Berger : Hasse’s class number product formula for generalized Dirichlet fields and other types of number fields, Manuscripta Math. 76 (1992), fasc. 3–4, 397–406; DOI: 10.1007/BF02567768 [y420155088532728.pdf (512.8 KB)];

PPN365956996_0076_log31.pdf (969,71 KB), GDZ; [PDF],

DigiZeitschriften; MR1185027 (93h:11121) (reviewed by S. Louboutin; reprinted in Reviews in Number Theory, volume 5B, R29-862) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0779.11051 (reviewed by M. Kolster) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[21] F. Lemmermeyer : Kuroda’s class number formula, Acta Arith. LXVI (1994), fasc. 3, 245–260; aa6633.pdf, icm; MR1276992 (95f:11090) (reviewed by C. J. Parry; reprinted in Reviews in Number Theory, volume 5B, R29-908) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0807.11052 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, Bib-TeX].

[22] P. J. Sime : On the ideal class group of real biquadratic fields, Trans. Amer. Math. Soc. 347 (1995), no. 12, 4855–4876; DOI: 10.1090/S0002-9947-1995-1333398-3 [PDF]; doi:10.2307/2155066 [PDF], JSTOR; MR1333398 (96c:11130) (reviewed by T. Metsänkylä; reprinted in Reviews in Number Theory, volume 5B, R29-930) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0847.11060 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[23] R. Kučera : On the parity of the class number of a biquadratic field, J. Number Theory 52 (1995), no. 1, 43– 52; doi:10.1006/jnth.1995.1054 [PDF (271 K)]; MR1331764 (96e:11139) (reviewed by J. Hurrelbrink; reprinted in Reviews in Number Theory, volume 5B, R29-935) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0852.11065 (reviewed by T. Lepistö) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[24] T. M. McCall, C. J. Parry, and R. R. Ranalli : Imaginary bicyclic biquadratic fields with cyclic 2-class group, J. Number Theory 53 (1995), no. 1, 88–99; doi:10.1006/jnth.1995.1079 [PDF (383 K)]; MR1344833 (96e:11131) (reviewed by B. de Smit; reprinted in Reviews in Number Theory, volume 5B, R16-708) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0831.11059 (reviewed by H. Yokoi) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[25] S.-I. Katayama and S. Katayama : Fibonacci, Lucas and Pell Numbers and class numbers of bicyclic biquadratic fields, Math. Japon. 42 (1995), no. 1, 121–126; MR1344638 (96f:11144) (reviewed by H. Yokoi; reprinted in Reviews in Number Theory, volume 5B, R29-937) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0830.11040 (reviewed by T. Lepistö) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[26] S. R. Ashford : Weil’s additive characters and class number parity, K-theory and algebraic geometry: connections with quadratic forms and division algebras Santa Barbara, CA, 1992), Proc. Sympos. Pure Math., 58, Part 2, Amer. Math. Soc., Providence, RI, 1995, pp. 51–61; MR1327288 (96c:11129) (reviewed by N. Yui) [html, PDF, DVI, PostScript, BibTeX];

[27] G. Hetrodt : Über einen Zusammenhang zwischen der Spektraltheorie automorpher Funktionen auf dem oberen Halbraum und den Klassenzahlen biquadratischer Zahlkörper, Schriftenreihe des Mathematischen Instituts der Universität Münster. 3. Serie, Heft 17, ii+80 S., Schriftenreihe Math. Inst. Univ. Münster 3. Ser., 17, Univ. Münster, Münster, 1996; MR1395381 (97j:11025) (reviewed by Y. Petridis) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0847.11030 (reviewed by J. Elstrodt) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[28] T. M. McCall, C. J. Parry, and R. R. Ranalli : The 2rank of the class group of imaginary bicyclic biquadratic fields, Canad. J. Math. 49 (1997), no. 2, 283–300; DOI: 10.4153/CJM-1997-014-2 [mccall0393.pdf [113 KB]]; MR1447492 (99d:11116) (reviewed by B. de Smit) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0884.11043 (reviewed by E. Benjamin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[29] A. J. Eichstädt : Automorphe Funktionen auf dem dreidimensionalen hyperbolischen Raum und Klassenzahlen biquadratischer Zahlkörper, Münster: Univ. Münster, Fachbereich Mathematik und Informatik, 2001, 82 S.; Zbl 1013.11020 (reviewed by J. A. Antoniadis) [html, PDF, XML, ASCII, DVI, PS, Bib-TeX].

[30] A. Azizi et A. Mouhib : Sur le rang du 2-groupe de classes de ℚ(\sqrt{m}, \sqrt{d}) où m = 2 où un premier p ≡ 1 (mod 4), Trans. Amer. Math. Soc. 353 (2001), no. 7, 2741–2752; DOI: 10.1090/S0002-9947(01)02753-2 [PDF]; doi:10.2307/221824 [PDF], JSTOR; MR1828471 (2002b:11152) (reviewed by A. Movahhedi) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0986.11073 (reviewed by E.Benjamin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[31] K. Wang : Fundamental unit system and class number of real bicyclic biquadratic number fields, Proc. Japan Acad. Ser. A Math. Sci. [日本学士院紀要] 77 (2001), No. 9, 147–150; doi:10.3792/pjaa.77.147, euclid.pja/1148393008; MR1869110 (2002j:11126) (reviewed by S. Louboutin) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1011.11071 (reviewed by H. Yokoi) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[32] A. Azizi et A. Mouhib : Le 2-rang du groupe de classes de certains corps biquadratiques et applications, Internat. J. Math. 15 (2004), no. 2, 169– 182; doi:10.1142/S0129167X04002235 [S0129167X04002235. pdf (234 KB)]; 12633348.pdf (243 KB), Academic Search Premier, EBSCO; MR2055368 (2005i:11154) (reviewed by J.-F. Jaulent) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1051.11057 (reviewed by E. Benjamin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[33] T. M. McCall, C. J. Parry, and R. R. Ranalli : Imaginary bicyclic biquadratic fields with elementary 2-class group, JP J. Algebra Number Theory Appl. 5 (2005), no. 1, 173–195; [PDF (187.8 K)]; MR2140320 (2006i:11124) (summary) [html, PDF, DVI, PostScript, BibTeX].

[34] T. M. McCall, C. J. Parry, and R. R. Ranalli : The structure of the 2-Sylow subgroups of the ideal class groups of imaginary bicyclic biquadratic fields, JP J. Algebra Number Theory Appl. 8 (2007), Issue 1, 81–105; [PDF (187.78 K)]; MR2370193 (2008m:11218) (reviewed by D. A. Buell) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1165.11079 (reviewed by P. Buckingham) [html, PDF, XML, ASCII, DVI, PS, Bib-TeX].

[35] 中島匠一 : 有理数体の (2, 2) 拡大の類数, 第 7 回北陸数論研究集会報告集 [Proceedings of the 7th Workshop on Number Theory] (金沢大学サテライトプラザ, 2008) (平林幹人, 野村明人, 山下浩, 木村巌, 菅野孝史編), 2009, pp. 46–58.

[36] 山村健 : (2, 2) 型の Abel (拡大) 体の類数および類群に関する文献 (2009 年 2 月更新版), 第 7 回北陸数論研究集会報告集 [Proceedings of the 7th Workshop on Number Theory] (金沢大学サテライトプラザ, 2008) (平林幹人, 野村明人, 山下浩, 木村巌, 菅野孝史編), 2009, pp. 134–138.